# Skyfall-GS: Synthesizing Immersive 3D Urban Scenes from Satellite Imagery

### v1

@@ A.2 Experiments
    Geometric Accuracy: "Which video’s 3D structures (buildings, terrain, objects) more accurately represent the real-world geometry when compared to the ground truth video?"
@@
    Spatial Alignment: "Which video’s layout and positioning of elements better matches the satellite imagery reference?"
@@
    Overall Perceptual Quality: "Considering all aspects (geometry, textures, lighting, consistency), which video presents a more convincing and high-quality 3D representation of the scene?"
For the user study on DFC2019 dataset, each participant viewed videos from Sat-NeRF (Marí et al., 2022), our method without IDU, and our complete method, alongside Google Earth Studio reference footage and the original satellite imagery. For the user study on the GoogleEarth dataset, each participant viewed videos from CityDreamer (Xie et al., 2024), GaussianCity (Xie et al., 2025b) and our complete method, alongside Google Earth Studio reference footage and the reference satellite imagery.
Comparison details. For quantitative comparisons with Sat-NeRF (Marí et al., 2022), Mip-Splatting (Yu et al., 2024) and our method without IDU refinement, we used consistent camera parameters across all methods: $17^{\circ}$ elevation angle, 328-unit radius, and $20^{\circ}$ field of view, with cameras targeting the AOI’s origin. For comparisons with CityDreamer (Xie et al., 2024) and GaussianCity (Xie et al., 2025b), we use $45^{\circ}$ elevation angle, 1067-unit radius, and $20^{\circ}$ field of view, with cameras also targeting the AOI’s origin. These parameters were selected to ensure equitable comparison with similar scene coverage across methods.
Additional comparison with CoR-GS (Zhang et al., 2024b). We additionally compare with the state-of-the-art sparse-view 3D reconstruction method, CoR-GS (Zhang et al., 2024b). We used the official CoR-GS codebase and extended its training to 30,000 iterations to match our own for a fair comparison. We use the same evaluation protocol described in the main paper, evaluating on the DFC 2019 and the GoogleEarth dataset (Xie et al., 2024). As shown in Table˜9 and Table˜8, our method consistently outperforms CoR-GS (Zhang et al., 2024b) across all metrics and scenes, validating our satellite-specific designs including opacity regularization and pseudo-depth supervision. Furthermore, our full method, which adds the synthesis stage driven by a curriculum-based iterative dataset update (IDU), provides a substantial additional improvement, highlighting the effectiveness of our complete pipeline.
Table 8: Quantitative comparison on each AOI of DFC2019 (Le Saux et al., 2019). Our method consistently outperforms baseline methods on distribution metrics and most pixel-level metrics, indicating superior image synthesis quality. Metrics are computed between renders from each method and reference frames from GES.
@@
JAX_004 Sat-NeRF 79.97 3.838 11.95 0.2290 0.8700
EOGS 107.23 5.913 8.22 0.1271 1.0174
CoR-GS 91.01 5.131 11.25 0.2554 0.9793
Ours 24.45 1.474 12.90 0.2446 0.846
JAX_068 Sat-NeRF 93.70 5.376 9.86 0.2607 0.8414
EOGS 85.57 5.516 6.39 0.1593 0.9953
CoR-GS 90.34 5.864 11.77 0.3230 1.0073
Ours 28.35 2.845 11.79 0.2931 0.8210
JAX_214 Sat-NeRF 90.76 5.376 8.97 0.2684 0.8394
EOGS 71.02 4.342 7.40 0.2293 0.8883
CoR-GS 86.33 5.258 11.66 0.4074 0.9079
Ours 26.69 1.964 12.24 0.3881 0.7420
JAX_260 Sat-NeRF 89.00 4.881 9.43 0.3172 0.9068
EOGS 87.15 5.372 7.04 0.1574 0.9342
CoR-GS 88.44 4.710 11.50 0.4162 0.8977
Ours 29.83 2.076 12.59 0.3574 0.7540
Table 9: Quantitative comparison with CityDreamer (Xie et al., 2024), GaussianCity (Xie et al., 2025b), CoR-GS (Zhang et al., 2024b) on each AOI of the GoogleEarth dataset (Xie et al., 2024). The results show that our approach consistently achieves the best performance, indicating superior geometric and perceptual fidelity compared to all baselines. Metrics are computed between renders from each method and reference frames from GES.
@@
004 CityDreamer 39.88 3.869 13.06 0.3519 0.5643
GaussianCity 28.71 2.710 14.00 0.3786 0.5656
CoR-GS 33.69 4.203 11.55 0.3440 0.6120
Ours 10.43 2.491 15.09 0.3793 0.3978
010 CityDreamer 34.29 4.270 12.24 0.1387 0.5544
GaussianCity 29.67 2.850 12.90 0.1661 0.5335
CoR-GS 29.75 3.672 12.90 0.1807 0.4209
Ours 11.03 1.631 13.58 0.1769 0.4073
219 CityDreamer 29.53 4.097 11.63 0.1344 0.5471
GaussianCity 23.72 3.224 12.37 0.1676 0.5254
CoR-GS 16.29 3.173 12.64 0.1792 0.3974
Ours 10.36 1.279 13.12 0.1699 0.3975
336 CityDreamer 42.38 4.372 13.39 0.4431 0.5654
GaussianCity 32.83 2.883 14.36 0.4533 0.5382
CoR-GS 29.55 3.958 14.29 0.4592 0.3879
Ours 7.83 2.635 15.32 0.4662 0.3719
Additional qualitative comparisons. Due to space constraints in the main paper, we present additional qualitative comparison results in this supplementary material for scenes JAX_004 and JAX_260 from the DFC2019 dataset, and scenes 004 and 219 from the GoogleEarth dataset. Figure˜11 shows orbital view comparisons with Sat-NeRF (Marí et al., 2022), Mip-Splatting (Yu et al., 2024), CoR-GS (Zhang et al., 2024b), and EOGS (Savant Aira et al., 2025), while Figure˜12 presents city-scale view comparisons with CityDreamer (Xie et al., 2024), GaussianCity (Xie et al., 2025b), and CoR-GS (Zhang et al., 2024b). These additional results further demonstrate the consistent superiority of our method across diverse urban environments.
Additional visual results. We also provide qualitative results on four additional AOIs from Jacksonville to demonstrate our method’s robustness across diverse urban environments. As shown in Figure˜13 and Figure˜14, these AOIs contain distinctive architectural features: JAX_004 showcases a residential area with mixed housing types and green spaces; JAX_164 features a prominent city hall building with its characteristic dome and symmetrical facade; JAX_175 encompasses an American football stadium with its distinctive oval structure and surrounding parking facilities; JAX_168 contains a commercial district with varied building heights and dense urban layout. Despite these varied urban typologies, our method successfully generates coherent three-dimensional renderings that preserve the spatial relationships and architectural features present in the satellite imagery. These additional results further validate the generalizability of our approach across diverse urban landscapes without requiring scene-specific parameter adjustments.
Multi-date Appearance Variation. The use of multi-date satellite imagery introduces a significant challenge, as images of the same location, when captured on different days, exhibit drastic variations in appearance. As showned in Figure 15, these differences can fundamentally alter the scene’s geometry and texture. Effectively synthesizing novel views requires a model capable of intelligently disentangling the static 3D scene structure from these challenging, temporally-varying appearance factors.
004 Refer to caption Refer to caption Refer to caption Refer to caption
068 Refer to caption Refer to caption Refer to caption Refer to caption
214 Refer to caption Refer to caption Refer to caption Refer to caption
260 Refer to caption Refer to caption Refer to caption Refer to caption
Scene Input Satellite Image 3DGS Render
@@
164 Refer to caption Refer to caption Refer to caption Refer to caption
168 Refer to caption Refer to caption Refer to caption Refer to caption
175 Refer to caption Refer to caption Refer to caption Refer to caption
264 Refer to caption Refer to caption Refer to caption Refer to caption
Scene Input Satellite Image 3DGS Render
@@
Figure 15: Visualization of multi-data satellite imagery of the DFC2019 dataset. Note the substantial shifts in appearance, including changes in illumination, cloud cover, and surface characteristics, which introduce challenges for consistent 3D reconstruction.
A.3 LLM Usage Disclosure
Large language models (LLMs) were used to assist in improving the clarity and conciseness of the writing and in searching for related work. All technical ideas, algorithm designs, experiments, and analyses were conceived, implemented, and validated by the authors. The authors have carefully verified all content and take full responsibility for the correctness and integrity of this paper.

### v2

@@ 0.C.2 Evaluation Protocol
    Geometric Accuracy: “Which video’s 3D structures (buildings, terrain, objects) more accurately represent the real-world geometry when compared to the ground truth video?”
@@
    Spatial Alignment: “Which video’s layout and positioning of elements better matches the satellite imagery reference?”
@@
    Overall Perceptual Quality: “Considering all aspects (geometry, textures, lighting, consistency), which video presents a more convincing and high-quality 3D representation of the scene?”
For the user study on DFC2019 dataset, each participant viewed videos from Sat-NeRF [56], EOGS [77], CoR-GS [114], Mip-Splatting [111], and our complete method, alongside Google Earth Studio reference footage and the original satellite imagery. For the user study on the GoogleEarth dataset, each participant viewed videos from CityDreamer [101], GaussianCity [103], CoR-GS [114], Mip-Splatting [111] and our complete method, alongside Google Earth Studio reference footage and the reference satellite imagery.
0.C.2.2 Comparison details.
For quantitative comparisons with Sat-NeRF [56], Mip-Splatting [111] and our method without IDU refinement, we used consistent camera parameters across all methods: $17^{\circ}$ elevation angle, 328-unit radius, and $20^{\circ}$ field of view, with cameras targeting the AOI’s origin. For comparisons with CityDreamer [101] and GaussianCity [103], we use $45^{\circ}$ elevation angle, 1067-unit radius, and $20^{\circ}$ field of view, with cameras also targeting the AOI’s origin. These parameters were selected to ensure equitable comparison with similar scene coverage across methods.
0.C.3 Results
0.C.3.1 Per-scene quantitative comparison.
We provide per-scene quantitative comparisons in Tables˜9 and 8. On the DFC2019 dataset (Table˜9), our method achieves the best $\text{FID}_{\text{CLIP}}$ and CMMD scores across all four AOIs by a substantial margin, reducing $\text{FID}_{\text{CLIP}}$ from the next-best $\sim$84 to $\sim$27 on average. Pixel-level metrics are more mixed: our method leads on PSNR in three out of four scenes, while CoR-GS occasionally achieves higher SSIM scores due to its tendency to produce overly smooth reconstructions, which artificially inflate SSIM as discussed in the main paper. LPIPS improvements are consistent, with our method achieving the best or second-best score on all four scenes. On the GoogleEarth dataset (Table˜8), our method achieves the best $\text{FID}_{\text{CLIP}}$ and CMMD in three out of four. The one exception is scene 219, a low-rise residential area with limited vertical facades, where Mip-Splatting achieves lower $\text{FID}_{\text{CLIP}}$ (7.06 vs. 7.80) and CMMD (1.589 vs. 2.640). We attribute this to this scene’s relatively flat geometry, which reduces the benefit of our curriculum-based facade synthesis and allows reconstruction-only baselines to perform competitively. This is consistent with our method’s design intent: the gains from diffusion-based refinement are most pronounced in scenes with significant occluded vertical structure. Despite this single exception, our method achieves the best average performance across both datasets, confirming its robustness across diverse urban typologies.
Table 8: Quantitative comparison on each AOI of DFC2019 [40]. Our method consistently outperforms baseline methods on distribution metrics and most pixel-level metrics, indicating superior image synthesis quality. Metrics are computed between renders from each method and reference frames from GES. Red, orange, and yellow indicate the best, second best, and third best results, respectively.
@@ 0.C.3 Results
[…]
EOGS 106.69 5.889 8.21 0.1266 1.0165
[…]
EOGS 86.87 5.378 7.04 0.1568 0.9319
[…]
Table 9: Quantitative comparison on each scenes of the GoogleEarth dataset [101]. The results show that our approach consistently achieves highly competitive performance, indicating superior geometric and perceptual fidelity across most metrics compared to the baselines. Metrics are computed between renders from each method and reference frames from GES. Red, orange, and yellow indicate the best, second best, and third best results, respectively.
@@
004 CityDreamer 34.54 4.297 13.06 0.3519 0.5643
[…]
010 CityDreamer 39.95 3.948 12.24 0.1387 0.5541
[…]
219 CityDreamer 42.54 4.444 11.63 0.1344 0.5465
[…]
336 CityDreamer 29.59 4.110 13.39 0.4430 0.5654
[…]
0.C.3.2 Additional qualitative comparisons.
Due to space constraints in the main paper, we present additional qualitative comparison results in this supplementary material for scenes JAX_004 and JAX_260 from the DFC2019 dataset, and scenes 004 and 219 from the GoogleEarth dataset. Figure˜12 shows orbital view comparisons with Sat-NeRF [56], Mip-Splatting [111], CoR-GS [114], and EOGS [77], while Figure˜13 presents city-scale view comparisons with CityDreamer [101], GaussianCity [103], and CoR-GS [114]. These additional results further demonstrate the consistent superiority of our method across diverse urban environments.
Figure 12: Additional qualitative comparison on the DFC2019 dataset with Sat-NeRF [56], Mip-Splatting [111], CoR-GS [114], and EOGS [77]. Our method significantly outperforms baseline approaches in both geometric accuracy and texture quality when rendering low-altitude novel views. Note the superior building geometry, facade details, and reduced floating artifacts in our final result.
Figure 13: Additional qualitative comparison on the GoogleEarth dataset with CityDreamer [101], GaussianCity [103], CoR-GS [114], and Mip-Splatting [111]. Our method is able to synthesize texture and geometry that is closer to the reference GES render.
Input Satellite Image 3DGS Render
004
Refer to caption Refer to caption Refer to caption Refer to caption
068
Refer to caption Refer to caption Refer to caption Refer to caption
214
Refer to caption Refer to caption Refer to caption Refer to caption
260
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Input Satellite Image 3DGS Render
164
Refer to caption Refer to caption Refer to caption Refer to caption
168
Refer to caption Refer to caption Refer to caption Refer to caption
175
Refer to caption Refer to caption Refer to caption Refer to caption
264
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 16: Visualization of multi-date satellite imagery of the DFC2019 dataset. Note the substantial shifts in appearance, including changes in illumination, transient objects, and surface characteristics, which introduce challenges for consistent 3D reconstruction.
0.C.3.3 Additional visual results.
We also provide qualitative results on four AOIs presents on the main paper (Table˜5), as well as four additional AOIs (Table˜6) from Jacksonville to demonstrate our method’s robustness across diverse urban environments. As shown in Figure˜14 and Figure˜15, these AOIs contain distinctive architectural features: JAX_004 showcases a residential area with mixed housing types and green spaces; JAX_164 features a prominent city hall building with its characteristic dome and symmetrical facade; JAX_175 encompasses an American football stadium with its distinctive oval structure and surrounding parking facilities; JAX_168 contains a commercial district with varied building heights and dense urban layout. Despite these varied urban typologies, our method successfully generates coherent three-dimensional renderings that preserve the spatial relationships and architectural features present in the satellite imagery. These additional results further validate the generalizability of our approach across diverse urban landscapes without requiring scene-specific parameter adjustments.
0.C.3.4 Multi-date appearance variation.
The use of multi-date satellite imagery introduces a significant challenge, as images of the same location, when captured on different days, exhibit drastic variations in appearance. As shown in Figure 16, these differences can fundamentally alter the scene’s geometry and texture. Effectively synthesizing novel views requires a model capable of intelligently disentangling the static 3D scene structure from these challenging, temporally-varying appearance factors.
Appendix 0.D Additional Experiments
0.D.0.1 Qualitative results on complex geometries.
To demonstrate the robustness of our framework beyond standard city-block layouts, we evaluate our method on scenes featuring irregular and historically significant architectures. As shown in Figure 15, we present synthesis results for Neuschwanstein Castle and Wells Cathedral. These scenes pose significant challenges due to their intricate non-Manhattan geometries, including sharp spires, varying elevations, and gothic architectural details. Despite these complexities, our method successfully disentangles the underlying geometry from the satellite input and hallucinates plausible high-frequency details for facades that are heavily occluded in the nadir views. This confirms that our hybrid reconstruction-generation approach is not limited to simple urban prisms but extends effectively to complex, free-form structures.
Input Satellite Image     3DGS Rendered Views
Neuschwanstein Castle
Refer to caption Refer to caption Refer to caption Refer to caption
Wells Cathedral
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 17: Qualitative results on complex geometries. Given multi-view satellite imagery, our method synthesizes novel views of irregular historical structures, shown here for Neuschwanstein Castle and Wells Cathedral.
0.D.0.2 Synthesis of bridges.
In addition to dense building clusters, we evaluate our method’s performance on scenes with complex topological structures, such as bridges. Figure 18 illustrates renders of bridges in JAX_068, JAX_214 and JAX_175, a typically difficult case for standard photogrammetry due to the thin structural components. Our method successfully recovers the connectivity of the bridge span while synthesizing realistic water textures. The diffusion-based refinement effectively regularizes the geometry, preventing the characteristic "melting" artifacts often observed in thin structures when using satellite-only reconstruction.
Refer to caption Refer to caption Refer to caption
(a) JAX_068 (b) JAX_175 (c) JAX_214
Figure 18: Qualitative results for bridges. We present the render results for bridges appears in JAX_068, JAX_214 and JAX_175, demonstrating the method’s ability to handle complex mulit-level structures that are typically challenging for standard reconstruction pipelines.
0.D.0.3 Visualizing transient object handling via per-image embeddings.
A key challenge in multi-date satellite reconstruction is the handling of dynamic elements, such as moving vehicles and pedestrians, which can introduce ghosting artifacts. Our approach addresses this by learning per-image appearance embeddings $e_{j}$ that capture photometric variations specific to each capture date. As visualized in Figure˜19, rendering the same viewpoint across 20 distinct appearance embeddings reveals that transient objects exhibit significant variability, appearing clearly in some embeddings while fading or vanishing in others. This qualitative evidence suggests that our appearance modeling effectively acts as a “sink” for transient data that does not align with the static 3D geometry. By absorbing these inconsistencies into the appearance code rather than the geometric parameters, the optimization naturally disentangles transient elements from the underlying static structure, ensuring a clean and consistent geometric reconstruction.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Emb. 00 Emb. 01 Emb. 02 Emb. 03 Emb. 04
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Emb. 05 Emb. 06 Emb. 07 Emb. 08 Emb. 09
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Emb. 10 Emb. 11 Emb. 12 Emb. 13 Emb. 14
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Emb. 15 Emb. 16 Emb. 17 Emb. 18 Emb. 19
Figure 19: Visualizing transient object handling via per-image embeddings. We render the same viewpoint using 20 different learned appearance embeddings (Emb. 00–19). Observe that transient objects, such as the vehicles on the road, exhibit varying degrees of visibility across different embeddings (e.g., clearly visible in some, faded or absent in others), while the static building geometry remains consistent. This qualitatively demonstrates that our per-image appearance modeling effectively disentangles transient elements from the underlying static 3D structure, preventing dynamic artifacts from corrupting the geometric reconstruction.
0.D.0.4 Sensitivity analysis of refinement text prompts.
A practical concern for deployment is whether our method is sensitive to the specific wording of the FlowEdit text prompts. To investigate this, we evaluate six prompting strategies listed in Table˜10, ranging from highly descriptive source/target pairs (Baseline) to vague descriptions (Vague Source, Vague Target), domain-specific variants that emphasize geometry or texture (Focus Geometry, Focus Texture), and a minimal context-free pair (Context Free). As shown in Figure˜20, the visual quality of the refined renders remains largely consistent across all strategies, with only minor differences in fine-grained texture details. This robustness suggests that our method does not require careful prompt engineering and is tolerant to moderate variations in prompt specificity.
Figure 20: Refine renders with different prompt strategies.
Table 10: List of text prompts used in sensitivity analysis. We evaluate six different prompting strategies to test the robustness of our method.
Strategy Source Prompt ($P_{\text{src}}$) Target Prompt ($P_{\text{tar}}$)
Baseline Satellite image of an urban area with modern and older buildings, roads, green spaces. Some areas appear distorted, with blurring and warping artifacts. Clear satellite image of an urban area with sharp buildings, smooth edges, natural lighting, and well-defined textures.
Vague Source A blurry satellite image of an urban area. Clear satellite image of an urban area with sharp buildings, smooth edges, natural lighting, and well-defined textures.
Vague Target Satellite image of an urban area with modern and older buildings, roads, green spaces. Some areas appear distorted, with blurring and warping artifacts. A clear satellite image of an urban area.
Focus Geometry Satellite image of an urban area with modern and older buildings, roads, green spaces. Some areas appear distorted, with blurring and warping artifacts. Clear satellite image of an urban area with geometrically precise buildings, flat rooftops, straight edges, and well-defined roads.
Focus Texture Satellite image of an urban area with modern and older buildings, roads, green spaces. Some areas appear distorted, with blurring and warping artifacts. Clear satellite image of an urban area with realistic, high-resolution textures, detailed facades, clear vegetation, and natural lighting.
No Prompt - -
Context Free distorted, blurring, warping artifacts clear, sharp, smooth edges, natural lighting, well-defined textures
0.D.0.5 Multi-block scalability via combined imagery.
To validate the scalability of our framework across larger, continuous urban regions, we combined satellite imagery from two adjacent areas (JAX_214 and JAX_260) into a single, unified dataset covering approximately 1km $\times$ 512m. To accommodate the expanded spatial extent, we scaled the Iterative Dataset Update (IDU) look-at grid to $6\times 3$, while maintaining our standard 5-episode refinement schedule (10,000 iterations per episode). The total training time for this expanded region took approximately 9 hours (1.5 hours for the initial reconstruction stage and 7.5 hours for the synthesis stage) on a single NVIDIA RTX A6000 GPU. The final converged scene yielded $\sim$3.5 million Gaussians with a peak final training memory footprint of $\sim$46GB. As shown in Figure˜21, the expanded area demonstrate seamless boundary consistency, particularly evident along the connecting highway structures, confirming that our method scales robustly to multi-block environments without introducing stitching artifacts at the region boundaries.
Figure 21: Multi-block reconstruction from combined JAX_214 and JAX_260 AOIs. Skyfall-GS produces a seamless, artifact-free 3D scene across a $\sim$1km $\times$ 512m urban region, demonstrating scalability to multi-block environments without boundary stitching artifacts.
0.D.0.6 Episode-vs-coverage analysis of curriculum strategy.
To quantify the effectiveness of the IDU module in revealing occluded regions, we present an Episode-vs-Coverage analysis (Figure 22). Since ground truth 3D geometry is unavailable for these satellite scenes, we use the final converged 3DGS model as a proxy for the total scene surface. We compute the cumulative coverage by optimizing a visibility attribute for every Gaussian point against the camera poses utilized in each episode. As shown in the figure, the coverage ratio steadily increases from $\sim$0.50 in Episode 1 to $\sim$0.75 in Episode 5. This consistent gain confirms that our curriculum strategy, which progressively lowers camera elevation from $85^{\circ}$ to $45^{\circ}$, successfully reveals and reconstructs vertical facade geometry that was initially occluded in the top-down satellite views. However, we acknowledge a limitation in this metric: because it calculates coverage based on reconstructed points, it cannot account for “true holes” (surface areas that were never generated at all because they were completely occluded from all sampled views). Future work could address this by dynamically sampling IDU cameras to target specific geometric uncertainties or detected holes.
Figure 22: Episode-vs-Coverage analysis. The plot illustrates the cumulative surface coverage ratio increasing across refinement episodes. The curriculum-based strategy effectively exposes occluded regions, particularly vertical facades, as the camera elevation descends.
0.D.0.7 Stochastic appearance diversity.
To demonstrate the generative capacity of our hybrid framework, we evaluate the stochastic diversity of the synthesized textures in Figure 23. By varying the random seed during the diffusion refinement stage while maintaining the same geometric initialization, our method produces diverse yet plausible surface details for identical underlying structures. As illustrated in the figure, detailed features such as the text on the red building signage vary distinctively (e.g., “Outeil” vs. “CUTAN”). Crucially, the macroscopic building footprint remains geometrically fixed, confirming that our framework successfully disentangles the reconstruction of physical geometry (grounded in satellite constraints) from the generative synthesis of high frequency appearance.
Figure 23: Demonstration of stochastic appearance diversity while preserving geometric consistency. Our method generates diverse plausible textures for identical underlying geometry across different random seeds. Notice how the red signage text on the building facade varies distinctively (e.g., “Outeil” vs. “CUTAN”) while the building’s structural footprint remains fixed, confirming that our framework successfully disentangles geometric reconstruction (grounded in satellite data) from generative appearance synthesis (variable via diffusion).